# A Comprehensive Survey on Self-Supervised Learning for Recommendation

### v2

@@ Abstract.
copyright: acmcopyrightccs: General and reference Surveys and overviewsccs: Information systems Recommender systems
@@ 1. Introduction
The field of recommender systems has undergone a revolutionary transformation, thanks to the strong representation capabilities of deep learning techniques. Neural network architectures have played a pivotal role in this paradigm shift. By harnessing the power of deep learning models, such as Recurrent Neural Networks (RNNs) (Jannach and Ludewig, 2017; Guo et al., 2020), Graph Neural Networks (GNNs) (Ying et al., 2018; Wang et al., 2019a), and Transformer architectures (Sun et al., 2019; Wu et al., 2020a), recommender systems have achieved an unprecedented level of understanding of user preferences and behaviors. Consequently, this deep understanding has paved the way for precise and personalized recommendations that cater to individual user needs and preferences.
Existing supervised learning methods heavily depend on having abundant labeled data for effective training. However, practical recommender systems often encounter the problem of data sparsity (Xia et al., 2022b). This means that real-life recommendation scenarios often suffer from a limited amount of available data or an insufficient number of labeled examples. Consequently, supervised learning methods encounter significant difficulties in effectively generalizing and accurately learning representations of user preferences in such scenarios. Fortunately, taking inspiration from the accomplishments of self-supervised learning (SSL) (Liu et al., 2021d), SSL techniques has proven to be beneficial for recommender systems in addressing the issue of data sparsity (Ren et al., 2023b; Xia et al., 2022b). Generally speaking, the key idea of SSL lies in leveraging the inherent structures or patterns within the data itself to create supervision signals for learning, without relying solely on externally labeled data. This allows recommender systems to utilize unlabeled data and extract meaningful representations, enabling them to make accurate predictions and recommendations even in the presence of data sparsity.
@@
Relationship with Previous Surveys. i) Extensive Collection of SSL Research. Our survey provides a significantly broader collection of self-supervised learning (SSL) works in the context of recommendation, encompassing over 170 papers, surpassing the previous surveys which covered approximately 80 papers (Yu et al., 2023b; Jing et al., 2023). In addition, we incorporate generative methods (e.g., mask autoencoding, denoised diffusion) and adversarial learning, which were not previously covered but have gained prominence according to (Liu et al., 2021d). ii) Comprehensive Taxonomy Design. Our taxonomy refines previous classifications of contrastive learning (Yu et al., 2023; Jing et al., 2023) and proposes a view-centric taxonomy approach. We also survey generative and adversarial learning methods based on their learning paradigms and targets. iii) Exploration of Distinct Scenarios. Distinguishing ourselves from earlier works (Yu et al., 2023b; Jing et al., 2023), which did not differentiate SSL-enhanced recommenders across different scenarios, we conducted an extensive exploration of nine scenarios individually. By surveying research works within diverse scenarios, we provide researchers with a more comprehensive understanding of the context-specific to each scenario and the corresponding challenges they present.
@@ 2. Preliminaries and Definition
In this section, we provide a concise introduction to the essential background knowledge relevant to our survey on self-supervised for learning recommendation. We begin by briefly outlining the preliminaries of the tasks in recommender systems. Subsequently, we describe the definition of self-supervised learning and introduce prominent learning paradigms: contrastive, generative, and adversarial approaches.
@@ 2.1. Recommender Systems
The research on recommendation encompasses a wide range of tasks in diverse scenarios, such as collaborative filtering, sequential recommendation, and multi-behavior recommendation. These tasks exhibit distinct data paradigms and objectives. Here, we begin by providing a general definition without delving into specific variations across diverse recommendation tasks. In recommender systems, there are two primary sets: the set of users, denoted as $\mathcal{U}=\{u_{1},...,u_{i},...,u_{|\mathcal{U}|}\}$, and the set of items, denoted as $\mathcal{V}=\{v_{1},...,v_{j},...,v_{|\mathcal{V}|}\}$. Then, an interaction matrix $\mathcal{A}\in\mathbb{R}^{|\mathcal{U}|\times|\mathcal{V}|}$ is utilized to represent the recorded interactions between users and items. In this matrix, an entry $\mathcal{A}_{i,j}$ is assigned a value of $1$ if user $u_{i}$ has interacted with item $v_{j}$, and 00 otherwise. The definition of interaction is adaptable to various scenarios and datasets (e.g., watching a movie, clicking or making purchases on an e-commerce website). Furthermore, in various recommendation tasks, there are distinct auxiliary observed data denoted as $\mathcal{X}$. For instance, in knowledge graph-enhanced recommendation, $\mathcal{X}$ incorporates a knowledge graph comprising external item attributes with different entity types and corresponding relationships. While in social recommendation, $\mathcal{X}$ encompasses user-level relationships, such as friendships. With the above definition, a recommendation model optimizes a prediction function $f(\cdot)$ that aims to accurately estimate the preference score between any user $u$ and item $v$:
(1) $y_{u,v}=f(\mathcal{A},\mathcal{X},u,v).$
The preference score $y_{u,v}$ represents the likelihood of user $u$ interacting with item $v$. Based on this score, the recommender system can provide recommendations of uninteracted items to each user by presenting a ranking list of items according to the estimated preference scores. In Section 4, we will delve into the data formulation of $(\mathcal{A},\mathcal{X})$ under different recommendation scenarios, complementing the demonstration of self-supervised learning in recommendation.
@@ 2.2. Self-supervised Learning in Recommendation
Over the past years, deep neural networks have demonstrated outstanding performance with supervised learning in various fields, including computer vision (Voulodimos et al., 2018), natural language processing (Deng and Liu, 2018) and recommender systems (Zhang et al., 2019). However, due to the heavy reliance on labeled data, supervised learning faces challenges when dealing with label sparsity, which is also a common issue in recommendation systems (Wu et al., 2021b; Yu et al., 2023b). To address this limitation, self-supervised learning has emerged as a promising approach, leveraging the data itself as labels for learning. In this section, we introduce three fundamental self-supervised learning methodologies: contrastive, generative, and adversarial.
@@ 2.2.1. Contrastive Learning
To fully leverage the inherent information within the data itself as supervision signals, contrastive learning has emerged as a prominent self-supervised learning approach (Jaiswal et al., 2020). The primary objective of contrastive learning is to maximize the agreement between different views augmented from the data. Formally, in contrastive learning for the recommendation, the objective is to minimize the following loss function (Wu et al., 2021a; Xie et al., 2022c):
@@
Here *$\mathcal{E}_{*}\circ\omega_{*}$ represents the view creation operations, which can vary across different contrastive learning-based recommenders. Each operation consists of data permutation processes, $\omega_{1}(\cdot)$ and $\omega_{2}(\cdot)$, which may involve dropping nodes/edges in graphs, as well as embedding encoding processes, $\mathcal{E}_{1}$ and $\mathcal{E}2$. The objective of minimizing $\mathcal{L}_{con}$ is to obtain robust encoding functions that maximize the agreement between the views. This cross-view agreement can be achieved through methods such as mutual information maximization or instance discrimination.
2.2.2. Generative Learning
Generative learning seeks to understand data structures and patterns to learn meaningful representations. It optimizes a deep encoder-decoder model that reconstructs missing or corrupted input data. The encoder, $\mathcal{E}(\cdot)$, creates latent representations from the input, while the decoder, $\mathcal{D}(\cdot)$, reconstructs the original data from the encoder output. The goal is to minimize the discrepancy between reconstructed and original data (Liu et al., 2022a) as follows:
@@ 2.2.2. Generative Learning
Here, $\omega$ represents an operation like masking or perturbation. $\mathcal{D}\circ\mathcal{E}$ denotes the process of encoding and decoding to reconstruct output. Recent studies have introduced a decoder-only architecture that effectively reconstructs data without an encoder-decoder setup. This approach uses a single model (e.g., Transformer (Vaswani et al., 2017)) for reconstruction, and is commonly applied in sequential recommendation with generative learning (Sun et al., 2019). The loss function $\mathcal{L}_{gen}$ format depends on the data type, with mean square loss for continuous data and cross-entropy loss for categorical data.
@@ 2.2.3. Adversarial Learning
Adversarial learning is a training method used to generate high-quality outputs, using a generator $\mathcal{G}(\cdot)$. What sets adversarial learning apart from generative learning is the inclusion of a discriminator $\Omega(\cdot)$, which determines whether a given sample is real or generated (Gui et al., 2021; Jabbar et al., 2021). In adversarial learning, the generator aims to enhance the quality of its generated outputs in order to deceive the discriminator. Consequently, the learning objective of adversarial learning can be defined as follows:
(4) $\underset{\mathcal{G}}{\min}\ \underset{\Omega}{\max}\ \{\mathbb{E}_{x\sim P(\mathcal{A},\mathcal{X})}[\log\ {\Omega}(x)]+\mathbb{E}_{\hat{x}\sim P(\mathcal{G}(\mathcal{A},\mathcal{X}))}[\log\ (1-{\Omega}(\hat{x}))]\}.$
Here, the variable $x$ represents a real sample obtained from the underlying data distribution, while $\hat{x}$ represents a synthetic sample generated by the generator $G(\cdot)$. During the training process, both the generator and discriminator enhance their capabilities through a competitive interplay. Ultimately, the generator strives to generate high-quality outputs that can be advantageous for downstream tasks. This training approach is commonly employed in various domains, including sequential recommendation, to enhance performance.
@@ 3. Taxonomy
In this section, we present our comprehensive taxonomy of self-supervised learning in recommender systems. As previously mentioned, the self-supervised learning paradigm can be categorized into contrastive learning, generative learning, and adversarial learning. Therefore, our taxonomy is built upon these three categories, providing more detailed insights into each. Our overall taxonomy is shown in Figure 1.
@@ 3.1. Contrastive Learning in Recommendation
The fundamental principle of contrastive learning (CL) involves maximizing the agreement between different views. Hence, we propose a view-centric taxonomy, which contains three key components to consider when applying contrastive learning: creating the views, pairing the views to maximize agreement, and optimizing the agreement.
@@ 3.1.1. View Creation
The created view emphasizes various data aspects for the model. It can incorporate global collaborative information to improve the recommender’s handling of global relationships (Xia et al., 2022b) or introduce random noise to enhance model robustness (Yu et al., 2022). We regard augmentation on input data (e.g., graphs, sequences, input features) (Yu et al., 2023b; Jing et al., 2023) as data-perspective view creation, and hidden feature augmentation during inference as feature-level view creation. Additionally, the model-based contrastive paradigm (Yu et al., 2023b) serves as model-level view generation. Thus, we propose a hierarchical taxonomy encompassing view-creation techniques from basic data-level to neural model-level.
  • Data-based view creation: In the realm of contrastive learning-based recommenders, diverse views are created by augmenting input data. These augmented data points are subsequently processed through deep neural recommenders. The resulting output embeddings from different views are then paired and utilized for contrastive learning. The augmentation methods for the original data vary depending on the recommendation scenario. For instance, graph-based data may employ techniques such as node/edge dropout or the addition of noisy edges (Wu et al., 2021b; Zhou et al., 2023a), while item sequences may make use of masking, cropping, and replacing (Sun et al., 2019; Zhou et al., 2020).
  • Feature-based view creation: In addition to generating views directly from the data, some methods consider conducting augmentation based on the hidden feature encoded during the models’ forward process. These representations can include node embedding during graph-based message passing or token vectors in the Transformer, for example. By applying various augmentation techniques or introducing random perturbations to the representations multiple times (Wu et al., 2021b; Yu et al., 2023a), the model’s final output can be considered as different views. One commonly used practice is to add random noise to the representations (Xie et al., 2021).
  • Model-based view creation: Both data-based and feature-based augmentation are non-adaptive since they are non-parametric. Consequently, various sophisticated methods have emerged to generate different views using neural modules. These views contain specific information based on the model design. For example, intent disentanglement neural modules can capture user intents (Ren et al., 2023b; Wu et al., 2023c), while hyperparagraph graph modules can capture global relationships (Xia et al., 2022b). In contrast, model-based view creation involves learnable parameters within the view generator that are adaptive to the learning objective and optimized during the learning process.
@@ 3.1.2. Pair Sampling
  • Natural Sampling: One common approach to pair sampling is straightforward and non-heuristic, which we refer to as natural sampling. Positive pairs are formed by different views generated from the same data sample, while negative pairs are formed by views from different data samples. In cases where a central view exists for all samples, such as a global view derived from the entire graph, the local-global relationship also forms positive pairs. This approach is widely applied in most contrastive learning recommenders.
  • Score-based Sampling: Another approach to pair sampling is score-based sampling. In this approach, a module calculates scores of pairs to determine positive or negative pairs. For instance, one relevant score is the distance between two views, which can be used to form pairs (Xia et al., 2022a; Qiu et al., 2022). Alternatively, clustering can be applied on views, where positive pairs are those within the same cluster, and negative pairs are those in different clusters (Nie et al., 2022).

### v3

@@ Abstract.
copyright: acmlicensedjournal: CSURjournalyear: 2025journalvolume: 1journalnumber: 1article: 1publicationmonth: 1doi: 10.1145/3746280ccs: General and reference Surveys and overviewsccs: Information systems Recommender systems
@@ 1. Introduction
Recommender systems have undergone revolutionary transformation through deep learning techniques. Neural architectures including RNNs (Jannach et al., 2017; Guo et al., 2020), GNNs (Ying et al., 2018; Wang et al., 2019c), and Transformers (Sun et al., 2019; Wu et al., 2020b) have enabled unprecedented understanding of user preferences, facilitating precise personalized recommendations. However, existing supervised learning methods require abundant labeled data, while practical recommender systems often face data sparsity (Xia et al., 2022a), limiting effective generalization and accurate preference learning. Inspired by self-supervised learning (SSL) successes (Liu et al., 2021a), SSL techniques have proven beneficial for addressing data sparsity in recommender systems (Ren et al., 2023a; Xia et al., 2022a). SSL leverages inherent data structures to create supervision signals without relying on external labels, enabling systems to utilize unlabeled data and extract meaningful representations for accurate predictions despite data sparsity.
@@
Relationship with Previous Surveys. i) Extensive Collection of SSL Research. Our survey provides a significantly broader collection of self-supervised learning (SSL) works in the context of recommendation, encompassing over 170 papers, surpassing the previous surveys which covered approximately 80 papers (Yu et al., 2023a; Jing et al., 2023). In addition, we incorporate reconstructive methods (e.g., mask autoencoding, denoised diffusion) and adversarial learning, which were not previously covered but have gained prominence according to (Liu et al., 2021a). ii) Comprehensive Taxonomy Design. Our taxonomy refines previous classifications of contrastive learning (Yu et al., 2023a; Jing et al., 2023) and proposes a view-centric taxonomy approach. We also survey reconstructive and adversarial learning methods based on their learning paradigms and targets. iii) Exploration of Distinct Scenarios. Distinguishing ourselves from earlier works (Yu et al., 2023a; Jing et al., 2023), which did not differentiate SSL-enhanced recommenders across different scenarios, we conducted an extensive exploration of nine scenarios individually. By surveying research works within diverse scenarios, we provide researchers with a more comprehensive understanding of the context-specific to each scenario and the corresponding challenges they present.
@@ 2. Preliminaries and Definition
In this section, we provide a brief introduction to the background relevant to our survey on self-supervised learning for recommendations. We start with an overview of the tasks in recommender systems and then define self-supervised learning, introducing key paradigms: contrastive, reconstructive, and adversarial approaches.
@@ 2.1. Recommender Systems
Recommendation research encompasses diverse tasks including collaborative filtering, sequential recommendation, and multi-behavior recommendation. We provide a general definition applicable across these scenarios. In recommender systems, there are two primary sets: users $\mathcal{U}=\{u_{1},...,u_{i},...,u_{|\mathcal{U}|}\}$ and items $\mathcal{V}=\{v_{1},...,v_{j},...,v_{|\mathcal{V}|}\}$. An interaction matrix $\mathcal{A}\in\mathbb{R}^{|\mathcal{U}|\times|\mathcal{V}|}$ represents recorded user-item interactions, where $\mathcal{A}_{i,j}=1$ if user $u_{i}$ interacted with item $v_{j}$, and 0 otherwise. Various recommendation tasks incorporate distinct auxiliary data $\mathcal{X}$ (e.g., knowledge graphs in KG-enhanced recommendation, user relationships in social recommendation). A recommendation model optimizes a prediction function $f(\cdot)$ that estimates preference scores: $y_{u,v}=f(\mathcal{A},\mathcal{X},u,v)$. The score $y_{u,v}$ indicates interaction likelihood, enabling item ranking and recommendation. In Section 4, we will explore the data formulation of $(\mathcal{A},\mathcal{X})$ under various recommendation scenarios, complementing the demonstration of self-supervised learning in recommendation.
@@ 2.2. Self-supervised Learning in Recommendation
Over the past years, deep neural networks have demonstrated outstanding performance with supervised learning in various fields, including computer vision (Voulodimos et al., 2018), natural language processing (Deng et al., 2018) and recommender systems (Zhang et al., 2019). However, due to the heavy reliance on labeled data, supervised learning faces challenges when dealing with label sparsity, which is also a common issue in recommendation systems (Wu et al., 2021a; Yu et al., 2023a). To address this limitation, self-supervised learning has emerged as a promising approach, leveraging the data itself as labels for learning. In this section, we introduce three fundamental self-supervised learning methodologies: contrastive, reconstructive, and adversarial.
@@ 2.2.1. Contrastive Learning
To fully leverage the inherent information within the data itself as supervision signals, contrastive learning has emerged as a prominent self-supervised learning approach (Jaiswal et al., 2020). The primary objective of contrastive learning is to maximize the agreement between different views augmented from the data. Formally, in contrastive learning for the recommendation, the objective is to minimize the following loss function (Wu et al., 2021b; Xie et al., 2022c):
@@
Here, $\mathcal{E}_{*}\circ\omega_{*}$ represents view creation operations that vary among different methods. Each operation includes data permutation processes, $\omega_{1}(\cdot)$ and $\omega_{2}(\cdot)$, which may involve dropping nodes/edges in graphs, along with embedding encoding processes, $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$. The goal of minimizing $\mathcal{L}_{con}$ is to derive robust encoding functions that maximize agreement between views, achieved through methods like mutual information maximization or instance discrimination.
2.2.2. Reconstructive Learning
Reconstructive learning seeks to understand data patterns to learn meaningful representations. It optimizes a deep encoder-decoder model that reconstructs missing or corrupted input data. The encoder, $\mathcal{E}(\cdot)$, creates latent representations from the input, while the decoder, $\mathcal{D}(\cdot)$, reconstructs the original data from the encoder output. The goal is to minimize the discrepancy between reconstructed and original data (Liu et al., 2022b) as follows:
@@ 2.2.2. Reconstructive Learning
Here, $\omega$ represents operations like masking. $\mathcal{D}\circ\mathcal{E}$ denotes the encoding and decoding process for output reconstruction. Recent studies introduce a decoder-only architecture that reconstructs data using a single model (e.g., Transformer (Vaswani et al., 2017)), commonly applied in sequential recommendation with reconstructive learning (Sun et al., 2019). The loss function $\mathcal{L}_{gen}$ varies by data type, using mean square loss for continuous data and cross-entropy loss for categorical data.
@@ 2.2.3. Adversarial Learning
Adversarial learning is a training method used to generate high-quality outputs, using a generator $\mathcal{G}(\cdot)$. What sets adversarial learning apart from reconstructive learning is the inclusion of a discriminator $\Omega(\cdot)$, which determines whether a given sample is real or generated (Gui et al., 2021; Jabbar et al., 2021). In adversarial learning, the generator aims to enhance the quality of its generated outputs in order to deceive the discriminator. Consequently, the learning objective of adversarial learning can be defined as follows:
[…]
Here, the variable $x$ represents a real sample from the data distribution, while $\hat{x}$ denotes a synthetic sample generated by the generator $G(\cdot)$. During training, both the generator and discriminator improve through competitive interplay. Ultimately, the generator aims to produce high-quality outputs beneficial for downstream tasks. This approach is commonly used across various domains, including sequential recommendation, to enhance performance.
2.2.4. Interpretability and Explainability
In recommender systems, interpretability (Wang et al., 2021b) and explainability (Zhang et al., 2020) have emerged as critical concepts for elucidating the mechanisms of black-box neural models. Interpretability emphasizes the intrinsic transparency of a recommender’s structure and decision-making processes, while explainability pertains to the post hoc generation of human-understandable rationales for the model’s recommendations. In the context of self-supervised learning, these techniques enhance our understanding of two essential attributes of modern recommenders: (i) For interpretability, some methods (Zhou et al., 2023a) construct structured causal graphs to facilitate an interpretable recommendation process and employ contrastive learning to mitigate exposure bias. Others (Wang et al., 2021b; Jiang et al., 2024b, a; Ren et al., 2023a) derive specific module architectures that inherently possess interpretability and utilize self-supervised learning to optimize model parameter training. (ii) Regarding explainability, prominent approaches (Zhuang et al., 2024; Liu et al., 2023c; Liao et al., 2025) leverage diverse contrastive learning techniques to improve the modeling and learning processes of explainable recommendations. Additionally, some methods (Wang et al., 2022b) utilize explanations generated by models for data augmentation in contrastive learning or develop multiple discriminators (Ren et al., 2020) in adversarial learning to elucidate recommendation behaviors.
2.2.5. Robustness against Noise
Self-supervised learning, originally developed to identify patterns in large-scale unlabeled datasets, inherently possesses the ability to mitigate the negative effects of noise, which is often unavoidable in such data. By integrating self-supervised learning techniques into recommender systems, we can enhance the robustness of models against both intentional and feature-specific noise present in recommendation data. For instance, in adversarial learning, several methods introduce adversarial perturbations (Yuan et al., 2019) to the original data, thereby training more robust models that can withstand noise. Additionally, in contrastive learning, training models using mutual information lower bounds (Yang et al., 2022; Xia et al., 2022a; Jiang et al., 2023) with positive-negative discrimination has demonstrated significant advantages in handling noisy data. Some methods (Yu et al., 2022, 2023b) also incorporate noise into the augmentation process to further improve training outcomes. Furthermore, in reconstructive learning, the task of reconstructing the original data patterns from noisy inputs inherently equips the model with resilience to noise. Representative methods in this domain include variational auto-encoders (Li et al., 2017) and diffusion-based approaches (Wang et al., 2023e; Jiang et al., 2024a).
@@ 3. Taxonomy
In this section, we present our comprehensive taxonomy of self-supervised learning in recommender systems. Different from previous works (Jing et al., 2023; Yu et al., 2023a) that either focus solely on contrastive learning or categorize SSL into contrastive, generative, predictive and hybrid methods, we categorize SSL into three main learning paradigms (i.e., contrastive, reconstructive and adversarial) and propose detailed taxonomies within each paradigm. Our framework is built upon these three categories, providing comprehensive insights into each approach’s methodologies and applications. Figure 1 illustrates this organizational structure and highlights the key distinctions between different SSL paradigms.
@@ 3.1. Contrastive Learning in Recommendation
The fundamental principle of contrastive learning (CL) involves maximizing the agreement between different views. Hence, we propose a view-centric taxonomy, as illustrated in Figure 2, which contains three key components to consider when applying CL: creating the views, pairing the views to maximize agreement, and optimizing the agreement.
@@ 3.1.1. View Creation
The created view emphasizes various data aspects for the model. It can incorporate global collaborative information to improve the recommender’s handling of global relationships (Xia et al., 2022a) or introduce random noise to enhance model robustness (Yu et al., 2022). We regard augmentation on input data (e.g., graphs, sequences, input features) (Yu et al., 2023a; Jing et al., 2023) as data-perspective view creation, and hidden feature augmentation during inference as feature-level view creation. Additionally, the model-based contrastive paradigm (Yu et al., 2023a) serves as model-level view generation. Thus, we propose a hierarchical taxonomy encompassing view-creation techniques from basic data-level to neural model-level.
  • Data-based view creation. In the realm of contrastive learning-based recommenders, diverse views are created by augmenting input data. These augmented data points are subsequently processed through deep neural recommenders. The resulting output embeddings from different views are then paired and utilized for contrastive learning. The augmentation methods for the original data vary depending on the recommendation scenario. For instance, graph-based data may employ techniques such as node/edge dropout or the addition of noisy edges (Wu et al., 2021a; Zhou et al., 2023c), while item sequences may make use of masking, cropping, and replacing (Sun et al., 2019; Zhou et al., 2020).
  • Feature-based view creation. In addition to generating views directly from the data, some methods consider conducting augmentation based on the hidden feature encoded during the models’ forward process. These representations can include node embedding during graph-based message passing or token vectors in the Transformer, for example. By applying various augmentation techniques or introducing random perturbations to the representations multiple times (Wu et al., 2021a; Yu et al., 2023b), the model’s final output can be considered as different views. One commonly used practice is to add random noise to the representations (Xie et al., 2021).
  • Model-based view creation. Both data-based and feature-based augmentation are non-adaptive since they are non-parametric. Consequently, various sophisticated methods have emerged to generate different views using neural modules. These views contain specific information based on the model design. For example, intent disentanglement neural modules can capture user intents (Ren et al., 2023a; Wu et al., 2023b), while hyperparagraph graph modules can capture global relationships (Xia et al., 2022a). In contrast, model-based view creation involves learnable parameters within the view generator that are adaptive to the learning objective and optimized during the learning process.
@@ 3.1.2. Pair Sampling
  • Natural Sampling. One common approach to pair sampling is straightforward and non-heuristic, which we refer to as natural sampling. Positive pairs are formed by different views generated from the same data sample, while negative pairs are formed by views from different data samples. In cases where a central view exists for all samples, such as a global view derived from the entire graph, the local-global relationship also forms positive pairs. This approach is widely applied in most contrastive learning recommenders.
  • Score-based Sampling. Another approach to pair sampling is score-based sampling. In this approach, a module calculates scores of pairs to determine positive or negative pairs. For instance, one relevant score is the distance between two views, which can be used to form pairs (Xia et al., 2022b; Qiu et al., 2022). Alternatively, clustering can be applied on views, where positive pairs are those within the same cluster, and negative pairs are those in different clusters (Nie et al., 2022).